Solve 5x + 9 - (2x-3)=2x-7
Answer by checkley77(12569)

5x+9-(2x-3)=2x-7
5x+9-2x+3=2x-7
5x-2x-2x=-7-9-3
x=-19 answer.
proof:
5*-19+9-(2*-19-3)=2*-19-7
-95+9+38+3=-38-7
-45=-45